OpenStudy (hanner_b_nanner):

I need some help please Fan&&Medal Can someone help me solve this using the Completing the Square method? I always get confused. h(x)= x^2 +6x +12

OpenStudy (johnweldon1993):

Sure :) \[\large x^2 + 6x + 12 = 0\]

OpenStudy (johnweldon1993):

Start by moving the constant to the other side of the equation subtract 12 from both sides... \[\large x^2 + 6x = -12\]

OpenStudy (johnweldon1993):

How...you take the coefficient of the 'x' term...which is 6 here... Divide it by 2....so we now have 3 And the square the result.....so 3^2 = 9 We add this 9 to both sides of the equation \[\large x^2 + 6x + 9 = -12 + 9\]

OpenStudy (johnweldon1993):

Simplify that right side a bit \[\large x^2 + 6x + 9 = -3\]

OpenStudy (johnweldon1993):

now we just write the left side as the sum of 2 squares \[\large (x + 3)^2 = -3\] Make sense there? we need 2 numbers that add to 6 but also multiply to make 9...so 3 and 3 instead of writing \[\large (x + 3)(x + 3) = -3\] we just write \[\large (x + 3)^2= -3\]

OpenStudy (hanner_b_nanner):

ok, that's what I got.. I just have a bad habit of putting the wrong sign in front of stuff and all that, lol :)) Thank you SO much!!!

OpenStudy (johnweldon1993):

No problem :) and if you keep going you will get an imaginary number answer Fun stuff :D lol :P

OpenStudy (hanner_b_nanner):

ughhh, lol I don't understand why you need to know about imaginary numbers... apparently they're not even real :P

OpenStudy (johnweldon1993):

....*ba dum *ching* .... :P lol :D

OpenStudy (johnweldon1993):

but you do know how to proceed right? Square root both sides \[\large (x + 3) = \pm \sqrt{-3}\] Subtract 3 from both sides and break it into the 2 equations \(\large x = i\sqrt{3} - 3\) and \(\large x = -i\sqrt{3} - 3\)

OpenStudy (hanner_b_nanner):

Thanks :D